OpenStudy (bugsbunny):

Find x to the nearest tenth. 6(6)=X(X+X-3)

OpenStudy (anonymous):

36 = x (2x - 3) 36 = 2x^2 - 3x 2x^2 - 3x -36 = 0

OpenStudy (bugsbunny):

Does it go any further than that? I got that far.

OpenStudy (hunus):

You have to use the quadratic formula to go farther than that.